Equations Reducible to Quadratic Form:
Use the given substitution to solve the equation:
x4 - 7x2 + 6 = 0 u=x2
I don't know how to enter the CARET symbol, but eh x is to the 4th power and the other x is to the 2nd power

Answer by mananth(16949)

x4 - 7x2 + 6 = 0
x^2=u
u^2-7u+6=0
u^2-6u-u+6=0
u(u-6)-1(u-6)=0
(u-6)(u-1)=0
u=6 OR 1
x^2= 6 OR x^2=1
x= +/-sqrt(6)
OR
x= +/-(1)
m.ananth@hotmail.ca